I need help with elimination.
x+2y=-1
2x-3y=12
Thank you!

Answer by mananth(16949)

1 x + 2 y = -1 .............1
2 x -3 y = 12 .............2
Eliminate y
multiply (1)by 3
Multiply (2) by 2
3 x 6 y = -3
4 x -6 y = 24
Add the two equations
7 x = 21
/ 7
x = 3
plug value of x in (1)
1 x + 2 y = -1
3 + 2 y = -1
2 y = -1 -3
2 y = -4
y = -2